OpenStudy (anonymous):

Please help me on these 2. Quickly :( NUMBER ONE..... 4x^5/6 in radical form? NUMBER TWO..... Rationalize the denominator of 7/(3+ SQRT 3)

OpenStudy (zehanz):

First. you've got to know this:\[a ^{\frac{ b }{ c }}=(\sqrt[c]{a})^b =\sqrt[c]{a^b}\]Then apply to 4x^(5/6):\[4x ^{\frac{ 5 }{ 6 }}=4\sqrt[6]{x^5}\]

OpenStudy (zehanz):

For the second one, you need to know a trick: multiply both nominator and denominator with 3 - sqrt 3:\[\frac{ 7 }{ 3+\sqrt{3} }\frac{ 3-\sqrt 3 }{ 3-\sqrt 3 }\] You'll remember that\[(a+b)(a-b)=a^2-b^2 \] Applying this gives:\[\frac{ 7(3-\sqrt 3) }{ 9 - 3 }=\frac{ 21-7\sqrt 3 }{ 6 }\] Once you've seen this trick, you will always remember it as a way to rationalize the denominator!